OpenStudy (camerondoherty):

Rhett decides to build a square room for his movie and music collection. If the area of the room is 9x2 − 6x + 1 square feet, what is the length of one side of the room? (3x + 1) feet (9x − 1) feet (3x − 1) feet (9x + 1) feet

OpenStudy (anonymous):

\[9x^2-6x+1\] is a perfect square what do you think it is the square of?

OpenStudy (anonymous):

just make a guess, i will tell you if you are wrong

OpenStudy (camerondoherty):

-6x?

OpenStudy (anonymous):

u need to find the two solutions to the equation

OpenStudy (anonymous):

lol no that is way wrong !

OpenStudy (camerondoherty):

lol u told me to take a guess xD

OpenStudy (camerondoherty):

its 9x right?

OpenStudy (anonymous):

\[9x^2-6x+1\] s the square of \(3x-1\) which is another way of saying \[(3x-1)^2=9x^2-6x+1\]

OpenStudy (anonymous):

(3x-1)(3x-1)

OpenStudy (anonymous):

when i said make a guess, i was hoping you would guess one of the possible answers given, not something else!

OpenStudy (camerondoherty):

xD

OpenStudy (anonymous):

answer is obviously the 3rd one

OpenStudy (camerondoherty):

this girl isn't tht bright... lol

OpenStudy (anonymous):

all of these problems have to do with recognizing what is a perfect square i don't actually believe that

OpenStudy (anonymous):

it is recognizing that \[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

in this case you have \[9x^2-6x+1\] which is the same as \[(3x-1)^2\] that is what you are being asked for

OpenStudy (camerondoherty):

ty!

OpenStudy (anonymous):

and what kind of name is "Rhett"? what is this, gone with the wind?

OpenStudy (anonymous):

yw

OpenStudy (camerondoherty):

Rhett?

OpenStudy (anonymous):

Lol gone with the wind

OpenStudy (camerondoherty):

lol